# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v1

@@ 9. Conclusion
In this work, we propose the general clone framework, which encompasses all decomposition methods, and identify the optimal bounds within the general clone framework. We also present an efficient algorithm for numerically computing these optimal bounds. With these results, we achieve the best-known bounds. Experiments demonstrate the tightness of our analysis. Additionally, we explore the joint composition of LDP protocols in the shuffle model for the first time.
We hope that this work contributes to both the practical deployment and the theoretical advancement of the shuffle model in differential privacy.
Appendix A Misunderstanding of the Privacy Blanket Framework
In this section, we discuss and correct a misunderstanding in the literature regarding the privacy blanket framework (Koskela et al., 2021).
The original paper on the privacy blanket framework demonstrates its usage with the example of $k$-Random Response (see Definition 5) (Balle et al., 2019). The privacy blanket distribution for $k$-RR is simply the uniform distribution over $[k]$. Let $N_{0}$ and $N_{1}$ denote the counts of $x_{1}^{0}$ and $x_{1}^{1}$ appearing in $Y_{|\boldsymbol{b}|}$ (see Definition 4), respectively. When $b_{1}=1$, it is easy to see that
$\Pr[P_{0}^{B}(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\mid b_{1}=0]=\Pr[P_{1}^{B}(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\mid b_{1}=0].$
When $b_{1}=0$, the first user commits their true value, i.e., $y_{1}^{b}=x_{1}^{b}$. In this case, using combinatorial analysis, we can derive the following (Balle et al., 2019):
$\forall Y_{|\boldsymbol{b}|}:\quad\frac{\Pr[P_{0}^{B}(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\mid b_{1}=0]}{\Pr[P_{1}^{B}(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\mid b_{1}=0]}=\frac{N_{0}}{N_{1}}.$
The mistake in (Koskela et al., 2021) lies in their characterization of the joint distribution of $N_{0}$ and $N_{1}$:
$N_{0}\sim\text{Bin}(n-1,\frac{1}{e^{\epsilon}+k-1})+1\text{ and }N_{1}\sim\text{Bin}(n-1,\frac{1}{e^{\epsilon}+k-1})$
under the condition $b_{1}=0$ and $(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\sim P_{0}^{B}$. In reality, $N_{0}$ and $N_{1}$ are not independent, as also noted in (Luo et al., 2022). The correct characterization is as follows: let $N\sim\text{Bin}(n-1,\frac{k}{e^{\epsilon_{0}}+k-1})$, then $N_{0}\sim\text{Bin}(N,\frac{1}{k})+1$, and $N_{1}\sim\text{Bin}(N-N_{0},\frac{1}{k-1})$.
This mistake affects both the theoretical analysis and the experimental results reported in (Koskela et al., 2021).
Appendix B Proof of Theorem 5

### v5

@@ 9 Conclusion
In this work, we propose the general clone framework, which encompasses all decomposition methods, and identify the optimal bounds within the general clone framework. We also present an efficient algorithm for numerically computing these optimal bounds. With these results, we achieve the best-known bounds. Experiments demonstrate the tightness of our analysis. Additionally, we present methods for computing optimal amplification bounds for both joint composition and parallel composition in the shuffle model. We hope that this work contributes to both the practical deployment and the theoretical advancement of the shuffle model in differential privacy.
Appendix A Optimal Decompositions for Common Local Randomizers
In this section, we derive the simplified optimal decomposition for several common local randomizers. Each of these randomizers satisfies the extreme property discussed in Section 4.4, and thus their simplified optimal decompositions consist of exactly five components, expressed as follows:
$\mathcal{R}(x_{1}^{0})$ $=e^{\varepsilon_{0}}p\mathcal{Q}_{e^{\varepsilon_{0}},1}+p\mathcal{Q}_{1,e^{\varepsilon_{0}}}+e^{\varepsilon_{0}}q\mathcal{Q}_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}+r\mathcal{Q}_{1,1},$
$\mathcal{R}(x_{1}^{1})$ $=p\mathcal{Q}_{e^{\varepsilon_{0}},1}+e^{\varepsilon_{0}}p\mathcal{Q}_{1,e^{\varepsilon_{0}}}+e^{\varepsilon_{0}}q\mathcal{Q}_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}+r\mathcal{Q}_{1,1},$
$\forall i\in[2,n],\ \mathcal{R}(x_{i})$ $=p\mathcal{Q}_{e^{\varepsilon_{0}},1}+p\mathcal{Q}_{1,e^{\varepsilon_{0}}}+q\mathcal{Q}_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}+r\mathcal{Q}_{1,1}+(1-2p-q-r)\mathcal{Q}_{i}.$
The values of the coefficients $p,q,r$ for several widely-used local randomizers are summarized in Table 2 and 3. We provide the details for computing the corresponding coefficients.
A.1 Binary local hash
Let $\mathcal{X}=[D]$. The Binary Local Hashing (BLH) mechanism is defined as $\mathsf{BLH}(x)=(h,\mathsf{RR}(h(x)))$, where $h:[D]\to[2]$ is a uniformly random function drawn from =$\mathcal{H}=\{f:[D]\to[2]\}$, and $\mathsf{RR}$ denotes the 2-ary randomized response mechanism (see Example 4.1 for the definition) [WBLJ17].
Define $h^{*}:[D]\to[2],\forall x,h^{*}(x)=1$ and $\bar{h}^{*}:[D]\to[2],\forall x,\bar{h}^{*}(x)=0$. When $h\in\{h^{*},\bar{h}^{*}\}$, the output distribution of $\mathsf{BLH}(x)$ is the same for all inputs $x$. We have
$\inf_{x}\Pr[\mathsf{BLH}(x)=(h^{*},0)]=\frac{1}{|\mathcal{H}|(1+e^{\varepsilon_{0}})},\quad\inf_{x}\Pr[\mathsf{BLH}(x)=(h^{*},1)]=\frac{e^{\varepsilon_{0}}}{|\mathcal{H}|(1+e^{\varepsilon_{0}})},$
$\inf_{x}\Pr[\mathsf{BLH}(x)=(\bar{h}^{*},0)]=\frac{e^{\varepsilon_{0}}}{|\mathcal{H}|(1+e^{\varepsilon_{0}})},\quad\inf_{x}\Pr[\mathsf{BLH}(x)=(\bar{h}^{*},1)]=\frac{1}{|\mathcal{H}|(1+e^{\varepsilon_{0}})}.$
When $h\notin\{h^{*},\bar{h}^{*}\}$, there must exist two inputs $a,b\in[D]$ such that $h(a)=0$ and $h(b)=1$. In this case, we have
$\inf_{x}\Pr[\mathsf{BLH}(x)=(h,0)]=\inf_{x}\Pr[\mathsf{BLH}(x)=(h,1)]=\frac{1}{|\mathcal{H}|(1+e^{\varepsilon_{0}})}.$
In summary, the blanket distribution of BLH is “nearly” uniform distribution over ×$\mathcal{H}\times[2]$:
$\omega(h,y)=\inf_{x}\Pr[\mathsf{BLH}(x)=(h,y)]/\gamma=\begin{cases}\frac{1}{\gamma(1+e^{\varepsilon_{0}})},&(h,y)\notin\{(h^{*},1),(\bar{h}^{*},0)\},\\ \frac{e^{\varepsilon_{0}}}{\gamma(1+e^{\varepsilon_{0}})},&(h,y)\in\{(h^{*},1),(\bar{h}^{*},0)\}.\end{cases}$
where $\gamma=\sum_{(h,y)}\inf_{x}\Pr[\mathsf{BLH}(x)=(h,y)]=\frac{2}{1+e^{\varepsilon_{0}}}+\frac{e^{\varepsilon_{0}}-1}{2^{D-1}(1+e^{\varepsilon_{0}})}$.
Recall the definition in Section 4.4:
$S_{r_{0},r_{1}}=\left\{y\ \middle|\ \frac{\mathcal{R}(x_{1}^{0})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{0},\frac{\mathcal{R}(x_{1}^{1})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{1}\right\}.$
It can be computed that
$S_{e^{\varepsilon_{0}},1}$ =$=\{(h,y)\mid h(x_{1}^{0})=y\wedge h(x_{1}^{1})\neq y\},$
$S_{1,e^{\varepsilon_{0}}}$ =$=\{(h,y)\mid h(x_{1}^{0})\neq y\wedge h(x_{1}^{1})=y\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ =$=\{(h,y)\mid h(x_{1}^{0})=y\wedge h(x_{1}^{1})=y\}-\{(h,y)\mid h=h^{*}\vee h=\bar{h}^{*}\},$
$S_{1,1}$ =$=\{(h,y)\mid h(x_{1}^{0})\neq y\wedge h(x_{1}^{1})\neq y\}\cup\{(h,y)\mid h=h^{*}\vee h=\bar{h}^{*}\}.$
Consequently, it follows that:
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)}-\frac{2}{|\mathcal{H}|(1+e^{\varepsilon_{0}})}$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)}-\frac{1}{2^{D-1}(1+e^{\varepsilon_{0}})},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{2e^{\varepsilon_{0}}}{|\mathcal{H}|(1+e^{\varepsilon_{0}})}$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{e^{\varepsilon_{0}}}{2^{D-1}(1+e^{\varepsilon_{0}})}.$
Table 2: Optimal decompositions for common local randomizers
($D$ denotes the size of the domain.) $p$ $q$ $r$ $k$-RR [War65] $\frac{1}{e^{\varepsilon_{0}}+k-1}$ 00 $\frac{k-2}{e^{\varepsilon_{0}}+k-1}$ BLH [WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}-\frac{1}{2^{D-1}(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{e^{\varepsilon_{0}}}{2^{D-1}(e^{\varepsilon_{0}}+1)}$ RAPPOR [EPK14] $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{e^{-\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}-\frac{e^{-\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{D}}$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}+\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{D}}$ OUE [WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{e^{-\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}-\frac{e^{-\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)^{D-1}}$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}+\frac{1}{2(e^{\varepsilon_{0}}+1)^{D-1}}$ HR [ASZ19] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$
Table 3: Optimal decompositions for common local randomizers (Asymptotic)
$p$ $q$ $r$
BLH ($D\gg 1$[WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$
RAPPOR ($D\gg 1$[EPK14] $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{1}{e^{\varepsilon_{0}/2}(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}$
OUE ($D\gg 1$[WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2e^{\varepsilon_{0}}(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}$
A.2 RAPPOR
RAPPOR first encodes the input $x\in[D]$ using unary encoding, yielding $\mathrm{UE}(x)\in\{0,1\}^{D}$, where $\forall i\in[D],\ \mathrm{UE}(x)[i]=\mathbbm{1}[x=i]$. It then applies $\frac{\varepsilon_{0}}{2}$-DP 2-ary randomized response independently to each bit of $\mathrm{UE}(x)$. Specifically, the output distribution of RAPPOR satisfies:
$\Pr[\mathrm{RAPPOR}(x)=\boldsymbol{y}]=\frac{e^{(D-d_{H}(\mathrm{UE}(x),\boldsymbol{y}))\cdot\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}},$
where $d_{H}(\mathrm{UE}(x),\boldsymbol{y})$ denotes the Hamming distance between two binary vectors in $\{0,1\}^{D}$ [EPK14].
As in the case of BLH, RAPPOR also exhibits an extreme case, namely $\boldsymbol{y}^{*}=(1,1,\dots,1)$. We have:
$\omega(\boldsymbol{y})=\inf_{x}\Pr[\mathrm{RAPPOR}(x)=\boldsymbol{y}]/\gamma=\begin{cases}\frac{e^{(D-1-d_{H}(\boldsymbol{y},\boldsymbol{0}))\cdot\varepsilon_{0}/2}}{\gamma(1+e^{\varepsilon_{0}/2})^{D}},&\boldsymbol{y}\neq\boldsymbol{y}^{*},\\ \frac{e^{\varepsilon_{0}/2}}{\gamma(1+e^{\varepsilon_{0}/2})^{D}},&\boldsymbol{y}=\boldsymbol{y}^{*}.\end{cases}$
where
$\gamma=\sum_{\boldsymbol{y}}\inf_{x}\Pr[\mathrm{RAPPOR}(x)=\boldsymbol{y}]$ $=\sum_{\boldsymbol{y}\in\{0,1\}^{D}}\frac{e^{(D-1-d_{H}(\boldsymbol{y},\boldsymbol{0}))\cdot\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}}+\frac{e^{\varepsilon_{0}/2}-e^{-\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}}$
$=\frac{1}{(1+e^{\varepsilon_{0}/2})^{D}}\sum_{i=0}^{D}\tbinom{D}{i}e^{(i-1)\cdot\varepsilon_{0}/2}+\frac{e^{\varepsilon_{0}/2}-e^{-\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}}$
$=e^{-\varepsilon_{0}/2}+\frac{e^{\varepsilon_{0}/2}-e^{-\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}}.$
According to the definition in Section 4.4:
$S_{e^{\varepsilon_{0}},1}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=0\},$
$S_{1,e^{\varepsilon_{0}}}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=1\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=1\}-\{\boldsymbol{y}^{*}\},$
$S_{1,1}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=0\}\cup\{\boldsymbol{y}^{*}\}.$
It can be computed that
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad$=$\quad$= $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad$=$\quad$= $\frac{e^{-\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}-\frac{e^{-\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad$=$\quad$= $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}+\frac{e^{\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}}.$
A.3 Optimal Unary Encode (OUE)
OUE (Optimized Unary Encoding) follows the same initial step as RAPPOR by applying unary encoding to the input $x$, resulting in $\mathrm{UE}(x)\in\{0,1\}^{D}$. However, it differs in how randomized response is applied: for each bit in $\mathrm{UE}(x)$, a 2-ary randomized response mechanism satisfying $\varepsilon_{0}$-DP is applied to the bits with value 0, while the bits with value 1 are replaced with uniformly random bits [WBLJ17].
The resulting probability distribution is given by:
$\Pr[\mathrm{OUE}(x)=\boldsymbol{y}]=\frac{e^{(D-1-d_{H}(\boldsymbol{y}_{-x},\boldsymbol{0}))\cdot\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})^{D-1}},$
where $d_{H}$ denotes the Hamming distance, and $\boldsymbol{y}_{-x}$ denotes the vector $\boldsymbol{y}$ with the $x$-th coordinate removed.
As in the case of RAPPOR, OUE also exhibits an extreme case, namely $\boldsymbol{y}^{*}=(1,1,\dots,1)$. We have:
$\omega(\boldsymbol{y})=\inf_{x}\Pr[\mathrm{OUE}(x)=\boldsymbol{y}]/\gamma=\begin{cases}\frac{e^{(D-1-d_{H}(\boldsymbol{y},\boldsymbol{0}))\cdot\varepsilon_{0}}}{2\gamma(1+e^{\varepsilon_{0}})^{D-1}},&\boldsymbol{y}\neq\boldsymbol{y}^{*},\\ \frac{1}{2\gamma(1+e^{\varepsilon_{0}})^{D-1}},&\boldsymbol{y}=\boldsymbol{y}^{*}.\end{cases}$
where
$\gamma=\sum_{\boldsymbol{y}}\inf_{x}\Pr[\mathrm{OUE}(x)=\boldsymbol{y}]$ $=\sum_{\boldsymbol{y}\in\{0,1\}^{D}}\frac{e^{(D-1-d_{H}(\boldsymbol{y},\boldsymbol{0}))\cdot\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})^{D-1}}+\frac{1-e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})^{D-1}}$
$=\frac{1}{2(1+e^{\varepsilon_{0}})^{D-1}}\sum_{i=0}^{D}\tbinom{D}{i}e^{(i-1)\cdot\varepsilon_{0}}+\frac{1-e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})^{D-1}}$
$=\frac{1+e^{\varepsilon_{0}}}{2e^{\varepsilon_{0}}}+\frac{1-e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})^{D-1}}.$
According to the definition in Section 4.4:
$S_{e^{\varepsilon_{0}},1}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=0\},$
$S_{1,e^{\varepsilon_{0}}}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=1\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=1\}-\{\boldsymbol{y}^{*}\},$
$S_{1,1}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=0\}\cup\{\boldsymbol{y}^{*}\}.$
It can be computed that
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad$=$\quad$= $\frac{1}{2(1+e^{\varepsilon_{0}})},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad$=$\quad$= $\frac{e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})}-\frac{e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})^{D-1}},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad$=$\quad$= $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}+\frac{1}{2(1+e^{\varepsilon_{0}})^{D-1}}.$
A.4 Hadamard Response
The Hadamard Response [ASZ19] assumes the input domain $\mathcal{X}=[D]$, where $D=2^{r}$ is a power of two. The mechanism maps each input $x\in[D]$ to a row of the normalized Hadamard matrix $H\in\{-1,1\}^{D\times D}$, and samples an output index $y\in[D]$ with probability proportional to $\exp\left(\frac{\varepsilon_{0}}{2}H_{x,y}\right)$, where $\varepsilon_{0}$ is the privacy budget. Each entry of the Hadamard matrix is defined as $H_{x,y}=(-1)^{x\odot y}$, where $x\odot y$ denotes the bitwise inner product modulo 2 between the binary representations of $x$ and $y$; that is, the sum (mod 2) of the component-wise products of their bits.
Formally, the Hadamard Response mechanism is defined as:
$\Pr[\mathrm{HR}(x)=y]=\frac{1}{Z}\exp\left(\frac{\varepsilon_{0}}{2}H_{x,y}\right),$
where $H\in\{-1,1\}^{D\times D}$ is the Hadamard matrix, and $Z$ is the normalization constant ensuring that the distribution sums to one.
By direct computation from the definition, the blanket distribution of the Hadamard Response is the uniform distribution over $[D]$. Its corresponding parameter is $\gamma=\frac{2}{1+e^{\varepsilon_{0}}}$.
According to the definition in Section 4.4:
$S_{e^{\varepsilon_{0}},1}$ =$=\{y|x_{1}^{0}\odot y=1\wedge x_{1}^{1}\odot y=0\},$
$S_{1,e^{\varepsilon_{0}}}$ =$=\{y|x_{1}^{0}\odot y=0\wedge x_{1}^{1}\odot y=1\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ =$=\{y|x_{1}^{0}\odot y=1\wedge x_{1}^{1}\odot y=1\},$
$S_{1,1}$ =$=\{y|x_{1}^{0}\odot y=0\wedge x_{1}^{1}\odot y=0\}.$
Due to the properties of the Hadamard matrix, we have $|S_{e^{\varepsilon_{0}},1}|=|S_{1,e^{\varepsilon_{0}}}|=|S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}|=|S_{1,1}|$. Therefore,
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad$=$\quad$= $\frac{1}{2(1+e^{\varepsilon_{0}})},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad$=$\quad$= $\frac{1}{2(1+e^{\varepsilon_{0}})},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad$=$\quad$= $\frac{1}{2(1+e^{\varepsilon_{0}})}.$
Appendix B GPARV for Laplace Mechanism
We consider the Laplace mechanism over the domain $\mathbb{X}=\{0,1\}$, defined by
$\mathcal{R}(x)=x+\text{Lap}\left(\frac{1}{\varepsilon_{0}}\right),$
where $\text{Lap}(\lambda)$ denotes the Laplace distribution with scale parameter $\lambda$. The corresponding distributions $\mathcal{R}(0)$, $\mathcal{R}(1)$, and the blanket distribution $\omega$ are given as follows:
$\mathcal{R}(0)(y)=\frac{\varepsilon_{0}}{2}\exp\left(-\varepsilon_{0}|y|\right),\quad\mathcal{R}(1)(y)=\frac{\varepsilon_{0}}{2}\exp\left(-\varepsilon_{0}|y-1|\right),$
$\omega(y)=\inf_{x\in\{0,1\}}\frac{\mathcal{R}(x)(y)}{\gamma}=\begin{cases}\frac{\varepsilon_{0}}{2\gamma}\exp\left(\varepsilon_{0}(y-1)\right),&y\leq 0.5,\\ \frac{\varepsilon_{0}}{2\gamma}\exp\left(-\varepsilon_{0}y\right),&y>0.5,\end{cases}$
where $\gamma=\exp(-\varepsilon_{0}/2)$.
The cumulative distribution function (CDF) of the privacy amplification random variable $L^{0,1}_{\varepsilon}$ is computed as:
$\Pr\left[L^{0,1}_{\varepsilon}\leq t\right]=\begin{cases}0,&t<\gamma(1-e^{\varepsilon_{0}+\varepsilon}),\\ \frac{1}{2}\sqrt{\frac{e^{\varepsilon}}{1-e^{\varepsilon_{0}/2}t}},&\gamma(1-e^{\varepsilon_{0}+\varepsilon})\leq t<\gamma(1-e^{\varepsilon}),\\ 1-\frac{1}{2}(e^{\varepsilon_{0}/2}t+e^{\varepsilon})^{-1/2},&\gamma(1-e^{\varepsilon})\leq t<\gamma(e^{\varepsilon_{0}}-e^{\varepsilon}),\\ 1,&t\geq\gamma(e^{\varepsilon_{0}}-e^{\varepsilon}).\end{cases}$
The CDF of $L^{0,1}_{\varepsilon}$ with parameters $\varepsilon_{0}=1.0$ and $\varepsilon=0.1$ is shown in Fig. 8.
Appendix C Proof of Theorem 12